How do you solve this:
A chemist has 12 liters of a 25% acid solution. How many liters of water must be added to obtain a 20% solution?

Answer by mananth(16946)

--------percent ---------------- quantity
Solution I 25 ---------------- 12 liters
Water 0 ---------------- x liters
Mixture 20 ---------------- 12 + x liters

25*12+0 x =20(12+x)
300= 240 + 20 x
-20 x = 240 - -300
-20 x = -60
/ -20
x = 3
3 liters of Water